OpenStudy (anonymous):

Find the indicated trigonometric value in the specified quadrant. Function: cot θ = -2 Quadrant: IV Trigonometric Value ; sin θ

OpenStudy (campbell_st):

draw a picture  find h using pythagoras's theorem h^2 = 1^1 + 2^2 \[h = \sqrt{5}\] 4th quadrant sin is negative then \[\sin \theta = \frac{-1}{\sqrt{5}}\]

OpenStudy (anonymous):

so i multiply top and bottom by √5. and get

OpenStudy (campbell_st):

oops cot = 1/tan so \[\tan \theta = -\frac{1}{2}\]

OpenStudy (anonymous):

i got -√5/5 as my final answer, does this look right?

OpenStudy (campbell_st):

if you want to rationalize the denominator the answer will be \[\sin \theta = -\frac{\sqrt{5}}{5}\]